# Energetic Resilience of Linear Driftless Systems

### v2

@@ Abstract
When a malfunction causes a control system to lose authority over a subset of its actuators, achieving a task may require spending additional energy in order to compensate for the effect of uncontrolled inputs. To understand this increase in energy, we introduce energetic resilience metrics that quantify the maximal additional energy required to achieve finite-time regulation in linear driftless systems that lose authority over some of their actuators. We first derive optimal control signals and minimum energies to achieve this task in both the nominal and malfunctioning systems. We then obtain a bound on the worst-case energy used by the malfunctioning system, and its exact expression in the special case of loss of authority over one actuator. Further considering this special case, we derive bounds on additive and multiplicative metrics for energetic resilience. A simulation example on a model of an underwater robot demonstrates that these bounds are useful in quantifying the increased energy used by a system suffering a partial loss of control authority.
I Introduction
@@
Control systems can suffer from a wide variety of failures, either due to system faults or adversarial attacks. Depending on their nature, failures can prevent a system from achieving a specific performance objective, including reachability or safety, even if mitigating actions are taken. Informally, a system is said to be resilient if its objectives can be achieved despite a failure. A well-publicized example of a failure was when the Nauka module docked to the International Space Station (ISS) in 2021. A software error resulted in an unexpected firing of the module’s thrusters, tilting the ISS from its nominal position for $47$ minutes by up to $540^{\circ}$, described as a “loss of attitude control” by the National Aeronautics and Space Administration (NASA) [1]. The failure mode in this event was that of partial loss of control authority, where the system lost authority over some of its thrusters, and other thrusters on the ISS had to be used to counteract the effect of the uncontrolled thrust.
@@
Motivated by this example, in this paper, we consider systems that are affected by complete loss in control authority over a subset of their actuators. In such a system, the uncontrolled inputs, potentially chosen by an adversary, can take on any values in the input space. However, these uncontrolled inputs are measurable, and can be used by the controlled inputs — under the authority of the system — in order to achieve a task. Such actuator faults are the most common cause of failure in spacecraft attitude control systems, accounting for around $44\%$ of all faults [2, 3]. While building redundancy in actuators [4, 5] or using control reconfiguration schemes [5, 6, 3] can aid resilience, this techniques can result in prohibitively high costs in control design.
Addressing this problem with classical control methods is bound to fail. Standard fault-tolerant control theories consider only limited actuator failure modes, such as actuators “locked” into producing constant inputs [7], or actuators with reduced effectiveness that remain controllable [8, 9]. Under a loss of control authority, uncontrolled inputs can take on values with the same magnitude as the controlled inputs, leading to the failure of robust control techniques [10] which require external disturbances to have much smaller magnitudes than the controlled inputs [11]. The uncontrolled inputs can also change significantly in a short period of time, leading to the failure of adaptive control strategies [11, 12].
@@
In [13], Bouvier and Ornik derived conditions for the resilient reachability of linear systems, i.e., reachability of a target set under any uncontrolled inputs. These conditions were used in [10] to design controllers that enable linear systems to achieve a target despite this malfunction. To quantify the additional time it may take for a target to be reached under a loss of control authority, Bouvier et al. introduced the notion of quantitative resilience in [14]. Quantitative resilience was defined as the maximal ratio of minimum reach times to achieve a target for a nominal system and a malfunctioning system. While initially defined for linear driftless systems in [14], this definition was extended to general linear systems in [15, 16].
@@
In this paper, we introduce metrics for a similar quantitative notion called energetic resilience. Instead of considering minimal reach times for the nominal and malfunctioning systems as in the works of [14] and [16], we consider the minimal control energies required to achieve drive the state to the origin from an initial condition, in a given finite time. In practical systems, the maximal additional control energy required to achieve this task is directly related to maximal consumption of additional resources, such as fuel in vehicular systems. Quantifying this is important in understanding the maximum capacity of resources required. Designing this resource capacity must take into account all possible malfunctioning inputs, which is not a straightforward problem. Considering minimal control energies instead of minimal reach times allows us to derive optimal control signals that achieve the task in both the nominal and malfunctioning driftless systems, which was not accomplished by [14, 10, 16]. We also derive a closed-form expression for the worst-case uncontrolled input that seeks to maximize the energy used by the malfunctioning system, which is not provided in the earlier works. Similar metrics were used to quantify the maximal cost of disturbance by [17], considering linear systems affected by external bounded disturbances with no bounds on the control inputs. In contrast to [17], in our setting of partial loss of control authority, the controlled inputs are bounded, and uncontrolled inputs can take on values with the same magnitude. As a preliminary step towards studying the energetic resilience of general linear and nonlinear systems, we focus on linear driftless systems in this paper. Despite their simplicity, driftless dynamics are sufficiently rich to characterize a variety of robotic and underwater systems [18, 19].
@@
The contributions of this paper are organized as follows. In Section II, we formulate the problem statement, providing the definitions of key quantities derived in this paper. Section III discusses the optimal control signals, minimal control energies and restrictions on the time required to achieve a task in both the nominal and malfunctioning systems. These results use a technical lemma that is proved using the calculus of variations. Further, we also consider the worst-case control energy for the malfunctioning system over all possible uncontrolled inputs, and derive an upper bound for this quantity. We conclude Section III by briefly considering the problem of final time design, i.e., choosing a task completion time that minimizes the control energy in the malfunctioning system. In Section IV, we consider the special case of losing authority over one actuator, deriving an exact expression for the worst-case control energy for the malfunctioning system, as well as the worst-case uncontrolled input achieving this energy. We also derive bounds on additive and multiplicative metrics for energetic resilience, both of which quantify how much additional control energy is used by the malfunctioning system compared to the nominal system. In Section V, we consider a model of an underwater robot and illustrate the applicability of our results. In particular, we present the two energetic resilience metrics as a function of distance of the initial condition from the origin, and show that these metrics accurately characterize the additional control energy used by the malfunctioning system.
I-A Notation and Facts
@@
The set $\mathbb{R}^{+}\coloneqq[0,\infty)$ is the set of all non-negative real numbers. For scalars $z\in\mathbb{R}$, define the $\operatorname*{sign}(\cdot)$ function as $\operatorname*{sign}(z)=z/|z|\in\{-1,+1\}$ if $z\neq 0$, with $\operatorname*{sign}(0)=0$. For vectors $z\in\mathbb{R}^{n}$, the $\operatorname*{sign}(\cdot)$ function operates elementwise on $z$. The $p$-norm of a vector $x\in\mathbb{R}^{n}$ is defined as $\|x\|_{p}\coloneqq\left(\sum_{i=1}^{n}|x_{i}|^{p}\right)^{1/p}$, with $\|x\|_{\infty}\coloneqq\max_{i}|x_{i}|$. For a matrix $L\in\mathbb{R}^{p\times q}$ with entries indexed $l_{ij}$, define the induced matrix norms $\|L\|_{1}\coloneqq\max_{j}\sum_{i=1}^{p}|l_{ij}|$ and $\|L\|_{\infty}\coloneqq\max_{i}\sum_{j=1}^{q}|l_{ij}|$. The Moore-Penrose inverse [20], also called the pseudoinverse of $L$, is denoted $L^{\dagger}$. For a continuous function $u:[0,t_{f}]\to\mathbb{R}^{p}$, the $\mathcal{L}_{2}$ norm is defined as $\|u\|_{\mathcal{L}_{2}}\coloneqq\sqrt{\int_{t=0}^{t_{f}}\|u(t)\|_{2}^{2}\leavevmode\nobreak\ \mathrm{d}t}$.
@@
II Problem Formulation
In this paper, we consider linear driftless systems of the form
$\dot{x}(t)=Bu(t),\leavevmode\nobreak\ \leavevmode\nobreak\ x(0)=x_{0}\neq 0,$ (1)
where $x(t)\in\mathbb{R}^{n}$ is the state, $u(t)\in\mathbb{R}^{m+p}$ is the control and $B\in\mathbb{R}^{n\times(m+p)}$ is the input matrix. The set of admissible controls $\operatorname*{\mathcal{U}}$ is defined as:
$\operatorname*{\mathcal{U}}\coloneqq\left\{u:\mathbb{R}^{+}\to\mathbb{R}^{m+p}:\|u(t)\|_{\infty}\leq 1\leavevmode\nobreak\ \text{ for all $t$}\right\},$ (2)
in line with prior work [16, 22]. We consider malfunctions that result in the system losing control authority over $p$ of its $m+p$ actuators. Then, the matrix $B$ and control $u(t)$ can be split into controlled and uncontrolled components, and the malfunctioning system can be written as:
@@ II Problem Formulation
where the subscripts $c$ and $uc$ denote controlled and uncontrolled respectively. Here, $B=[B_{c},B_{uc}]$, $B_{c}\in\mathbb{R}^{n\times m}$, $B_{uc}\in\mathbb{R}^{n\times p}$, $u_{c}(t)\in\mathbb{R}^{m}$ and $u_{uc}(t)\in\mathbb{R}^{p}$. Based on the set of admissible controls $\operatorname*{\mathcal{U}}$ in (2), we have $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ and $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$, where
$\operatorname*{\mathcal{U}_{c}}$$\coloneqq\left\{u_{c}:\mathbb{R}^{+}\to\mathbb{R}^{m}:\|u_{c}(t)\|_{\infty}\leq 1\leavevmode\nobreak\ \text{for all $t$}\right\},$ (4a)
$\operatorname*{\mathcal{U}_{uc}}$$\coloneqq\left\{u_{uc}:\mathbb{R}^{+}\to\mathbb{R}^{p}:\|u_{uc}(t)\|_{\infty}\leq 1\leavevmode\nobreak\ \text{for all $t$}\right\}.$ (4b)
In this setting, the system has authority over the controlled input $u_{c}$, but the uncontrolled input $u_{uc}$ can be chosen arbitrarily from the space $\operatorname*{\mathcal{U}_{uc}}$, potentially by an adversary. However, this uncontrolled input is observable, and hence can be used in the design of $u_{c}$.
Our objective is the task of finite-time regulation, achieving $x(t_{f})=0$ for a specified final time $t_{f}$. In the nominal case, this task is achieved using the input $u\in\operatorname*{\mathcal{U}}$, while in the malfunctioning case, this task must be achieved using only the controlled input $u_{c}\in\operatorname*{\mathcal{U}_{c}}$, for any given uncontrolled inputs $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$. In this context, we define the notion of finite-time stabilizing resilience of a system, adapted from [14].
Definition 1 (Finite-time Stabilizing Resilience).
A system (1) is resilient to the loss of control authority over $p$ of its actuators, represented by the matrix $B_{uc}$ if, for all uncontrolled inputs $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$ and for a given final time $t_{f}$, there exists a controlled input $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ such that the system achieves the objective of finite-time regulation, i.e. $x(t_{f})=0$.
A system may be finite-time stabilizing resilient for some final times $t_{f}$ and not resilient for other final times. Throughout the rest of this paper, we refer to finite-time stabilizing resilience as simply resilience. We also assume that the nominal dynamics (1) are controllable, i.e., that finite-time regulation can be achieved from any initial state $x_{0}$. In both the nominal and malfunctioning cases, we are interested in the minimal control energies to achieve this task. Achieving this task might require considerably more control energy in the malfunctioning case compared to the nominal case. We thus aim to quantify the maximal additional control energy used by the malfunctioning system over all possible uncontrolled inputs $u_{uc}$, compared to the nominal system. To this end, we make the following definitions.
Definition 2 (Nominal Energy).
@@
Definition 3 (Malfunctioning Energy).
@@
Definition 4 (Total Energy).
@@
Definition 5 (Worst-case Total Energy).
The worst-case total energy is the maximal total energy (7) over all possible uncontrolled inputs $u_{uc}$:
$\overline{E}_{M}(x_{0},t_{f})$ $=\sup_{u_{uc}\in\operatorname*{\mathcal{U}_{uc}}}E_{M}^{+}(x_{0},t_{f},u_{uc})$
$=\sup_{u_{uc}\in\operatorname*{\mathcal{U}_{uc}}}\left\{E_{M}^{*}(x_{0},t_{f},u_{uc})+\left\|u_{uc}\right\|_{\mathcal{L}_{2}}^{2}\right\}$ (8)
The worst-case total energy quantifies the maximal effect of the uncontrolled input $u_{uc}$ on the overall control energy used by the system.
The overarching aim of this paper is to understand how much larger the worst-case total energy is, compared to the nominal energy. This is quantified by the following energetic resilience metrics, similar to those defined in [17].
Definition 6 (Additive Energetic Resilience).
For an initial condition $x_{0}$ no farther than a distance of $R$ from the origin, we define the additive energetic resilience of system (1) as
$r_{A}(t_{f},R)\coloneqq\sup_{\|x_{0}\|_{2}\leq R}\left\{\overline{E}_{M}(x_{0},t_{f})-E_{N}^{*}(x_{0},t_{f})\right\}.$ (9)
Definition 7 (Multiplicative Energetic Resilience).
For an initial condition at a distance of at least $R$ from the origin, we define the multiplicative energetic resilience of system (1) as
@@
Using (9), the malfunctioning system will use at most $r_{A}(t_{f},R)$ more energy than the nominal system to achieve finite-time regulation from an initial condition $x_{0}$ no farther than $R$ from the origin, and for a given $t_{f}$. We are thus interested in an upper bound for this metric. The constraint $\|x_{0}\|_{2}\leq R$ is required to ensure that this metric takes on a finite value. Similarly, (10) is a multiplicative measure of the increased energy required. For instance, if $r_{A}(t_{f},R)\geq 1/2$, then the actuators of the malfunctioning system use at most twice the energy used by the actuators of the nominal system to achieve finite-time regulation, for a given $t_{f}$ and $R$. We are thus interested in a lower bound for this metric. Without the constraint $\|x_{0}\|_{2}\geq R$, $x_{0}$ can be arbitrarily close to the origin and so can $E_{N}^{*}(x_{0},t_{f})$, reducing the metric to 00. We note that the definition of $r_{M}(t_{f},R)$ is closely related to the definition of quantitative resilience defined by Bouvier et al. for driftless systems in [14] and for general linear systems in [16]. However, quantitative resilience was defined based on reachable time while our definitions consider the control energy.
In the following section, we derive closed-form expressions for the nominal and malfunctioning energy, as well as a bound on the worst-case total energy. To quantify the maximal additional energy used by the malfunctioning system compared to the nominal system, we compare the quantities $E_{N}^{*}(x_{0},t_{f})$ from (5) and $\overline{E}_{M}(x_{0},t_{f})$ from (8). This comparison is achieved in Section IV by deriving bounds on the energetic resilience metrics (9) and (10), for the special case of losing authority over a single actuator.
III Nominal and Malfunctioning Energies
In this section, we derive expressions for the three quantities defined in Section II, as well as the corresponding optimal control inputs. We also briefly discuss a problem of final time design for the malfunctioning system, which finds the optimal final time that minimizes the malfunctioning energy for a given uncontrolled input.
Next, we present a lemma which is central to the development in this section.
Lemma 1.
Let $\mathcal{Z}$ be a set of continuous, real vector-valued functions with a given mean value $\overline{z}$ in the interval $[0,t_{f}]$, i.e.
$\mathcal{Z}\coloneqq\left\{z:[0,t_{f}]\to\mathbb{R}^{n_{z}}:\frac{1}{t_{f}}\int_{0}^{t_{f}}z(t)\mathrm{d}t=\overline{z}\in\mathbb{R}^{n_{z}}\right\}.$
Let $z^{*}(t)$ denote the signal with the minimum energy in $\mathcal{Z}$, i.e. $z^{*}\coloneqq\operatorname*{arg\,min}_{z\in\mathcal{Z}}\int_{0}^{t_{f}}z^{T}(t)z(t)\mathrm{d}t.$ Then,
$z^{*}(t)=\overline{z}\leavevmode\nobreak\ \text{ for all $t\in[0,t_{f}]$.}$
In other words, over all signals with a given mean value, the signal with the minimum energy is a constant equal to that mean value for all time.

### v3

@@ Abstract
When a malfunction causes a control system to lose authority over a subset of its actuators, achieving a task may require spending additional energy in order to compensate for the effect of uncontrolled inputs. To understand this increase in energy, we introduce an energetic resilience metric that quantifies the maximal additional energy required to achieve finite-time regulation in linear driftless systems that suffer this malfunction. We first derive optimal control signals and minimum energies to achieve this task in both the nominal and malfunctioning systems. We then obtain a bound on the worst-case energy used by the malfunctioning system, and its exact expression in the special case of loss of authority over one actuator. Further considering this special case, we derive a bound on the metric for energetic resilience. A simulation example on a model of an underwater robot demonstrates that this bound is useful in quantifying the increased energy used by a system suffering such a malfunction.
keywords:
Driftless systems, linear systems, optimal control, optimization, resilient control.
thanks: This work was supported by Air Force Office of Scientific Research grant FA9550-23-1-0131 and NASA University Leadership Initiative grant 80NSSC22M0070. (Corresponding Author: Ram Padmanabhan. Email: ramp3@illinois.edu.).
1 Introduction
@@
Control systems can suffer from a variety of failures, either due to system faults or adversarial attacks. Such failures can prevent a system from achieving a specific performance objective, including reachability or safety. Informally, a system is said to be resilient if its objectives can be achieved despite a failure. An example of a failure was the “loss of attitude control” (Bartels, 2021) suffered by the Nauka module while docking to the International Space Station (ISS) in 2021. The failure mode in this event was that of partial loss of control authority, where the module lost authority over some of its thrusters, and other thrusters on the ISS had to be used to counteract this uncontrolled thrust.
@@
Motivated by this example, in this paper, we consider systems that are affected by complete loss in control authority over a subset of their actuators. In such a system, the uncontrolled inputs, potentially chosen by an adversary, can take on any values in the input space. However, these uncontrolled inputs are measurable, and can be used by the controlled inputs — under the authority of the system — in order to achieve a task. We note that building actuator redundancy (Suich and Patterson, 1990; Grossman, 1995) or using control reconfiguration schemes (Boche et al., 2019; Hu et al., 2021) can aid resilience, but result in prohibitively high costs in control design. Moreover, standard fault-tolerant control considers only limited actuator failure modes (Tao et al., 2002; Amin and Hasan, 2019), and adaptive and robust control strategies fail since the uncontrolled input is unrestricted (Bouvier and Ornik, 2022; Anderson and Dehghani, 2008).
@@
Bouvier and Ornik (2020) derived conditions for the resilient reachability of linear systems, i.e., reachability of a target set under any uncontrolled inputs. These conditions were used by Bouvier and Ornik (2022) to design controllers that enable linear systems to achieve a target despite this malfunction. To quantify the additional time it may take for a target to be reached under a loss of control authority, Bouvier et al. (2021); Bouvier and Ornik (2023) introduced the notion of quantitative resilience for linear driftless and general linear systems. Quantitative resilience was defined as the maximal ratio of minimum reach times to achieve a target between a nominal system and a malfunctioning system.
@@
In this paper, we introduce a metric for a similar quantitative notion called energetic resilience. Instead of considering minimal reach times for the nominal and malfunctioning systems as in the works of Bouvier et al. (2021) and Bouvier and Ornik (2023), we consider the minimal control energies required to drive the state to the origin from an initial condition, in a given finite time. In practical systems, the maximal additional control energy required to achieve this task is directly related to maximal additional resource consumption, such as fuel in vehicles. Designing this resource capacity must take into account all possible malfunctioning inputs, which is not a straightforward problem. Considering minimal control energies instead of minimal reach times allows us to derive optimal control signals that achieve the task in both the nominal and malfunctioning driftless systems, which was not accomplished by Bouvier et al. (2021); Bouvier and Ornik (2022, 2023). We also derive a closed-form expression for the worst-case uncontrolled input that seeks to maximize the energy used by the malfunctioning system, which is not provided in the earlier works. A similar metric was used to quantify the maximal cost of disturbance by Padmanabhan et al. (2024), considering linear systems affected by external bounded disturbances with no input constraints. In contrast, this paper considers bounded controlled inputs and uncontrolled inputs can take on values with the same magnitude. As a preliminary step towards studying the resilience of general linear and nonlinear systems, we focus on linear driftless systems, which can characterize a variety of robotic and underwater systems Yu et al. (2016).
@@
Our contributions are organized as follows. In Section 2, we formulate the problem statement, providing the definitions of key quantities derived in this paper. Section 3 discusses the optimal control signals, minimal control energies and restrictions on the time required to achieve a task in both the nominal and malfunctioning systems. These results use a technical lemma characterizing the optimal control. We also derive an upper bound on the worst-case control energy for the malfunctioning system over all possible uncontrolled inputs. In Section 4, we consider the special case of losing authority over one actuator, deriving an exact expression for the worst-case control energy for the malfunctioning system and the worst-case uncontrolled input achieving this energy. We also derive a bound on the metric for energetic resilience, quantifying how much additional control energy is used by the malfunctioning system compared to the nominal system. In Section 5, we consider a model of an underwater robot and illustrate the applicability of our results. In particular, we show that the resilience metric accurately characterizes the additional control energy used by the malfunctioning system.
1.1 Notation and Facts
@@
The set $\mathbb{R}^{+}\coloneqq[0,\infty)$ is the set of all non-negative real numbers. For scalars $z\in\mathbb{R}$, define the $\operatorname*{sign}(\cdot)$ function as $\operatorname*{sign}(z)=z/|z|\in\{-1,+1\}$ if $z\neq 0$, with $\operatorname*{sign}(0)=0$. The $\operatorname*{sign}(\cdot)$ function operates elementwise on vectors $z\in\mathbb{R}^{n}$. The $p$-norm of a vector $x\in\mathbb{R}^{n}$ is defined as $\|x\|_{p}\coloneqq\left(\sum_{i=1}^{n}|x_{i}|^{p}\right)^{1/p}$, with $\|x\|_{\infty}\coloneqq\max_{i}|x_{i}|$. For a matrix $L\in\mathbb{R}^{p\times q}$ with entries indexed $l_{ij}$, define the induced matrix norms $\|L\|_{1}\coloneqq\max_{j}\sum_{i=1}^{p}|l_{ij}|$ and $\|L\|_{\infty}\coloneqq\max_{i}\sum_{j=1}^{q}|l_{ij}|$. The Moore-Penrose inverse (Penrose, 1955), also called the pseudoinverse of $L$, is denoted $L^{\dagger}$. For a continuous function $u:[0,t_{f}]\to\mathbb{R}^{p}$, the $\mathcal{L}_{2}$ norm is defined as $\|u\|_{\mathcal{L}_{2}}\coloneqq\sqrt{\int_{t=0}^{t_{f}}\|u(t)\|_{2}^{2}~{}\mathrm{d}t}$.
@@
2 Problem Formulation
In this paper, we consider linear driftless systems
$\dot{x}(t)=Bu(t),~{}~{}x(0)=x_{0}\neq 0,$ (1)
where $x(t)\in\mathbb{R}^{n}$ is the state and $u(t)\in\mathbb{R}^{m+p}$ is the control. The set of admissible controls $\operatorname*{\mathcal{U}}$ is defined as:
$\operatorname*{\mathcal{U}}\coloneqq\left\{u:\mathbb{R}^{+}\to\mathbb{R}^{m+p}:\|u(t)\|_{\infty}\leq 1~{}\text{ for all $t$}\right\},$ (2)
in line with prior work (Bouvier and Ornik, 2023). We consider malfunctions that result in the system losing control authority over $p$ of its $m+p$ actuators. Then, the matrix $B$ and control $u(t)$ can be split into controlled and uncontrolled components:
@@ 2 Problem Formulation
where the subscripts $c$ and $uc$ denote controlled and uncontrolled respectively. Here, $B=[B_{c},B_{uc}]$, $B_{c}\in\mathbb{R}^{n\times m}$, $B_{uc}\in\mathbb{R}^{n\times p}$, $u_{c}(t)\in\mathbb{R}^{m}$ and $u_{uc}(t)\in\mathbb{R}^{p}$. The set of admissible controls $\operatorname*{\mathcal{U}}$ splits into $\operatorname*{\mathcal{U}_{c}}$ and $\operatorname*{\mathcal{U}_{uc}}$ as
$\operatorname*{\mathcal{U}_{c}}$$\coloneqq\left\{u_{c}:\mathbb{R}^{+}\to\mathbb{R}^{m}:\|u_{c}(t)\|_{\infty}\leq 1~{}\text{for all $t$}\right\},$ (4a)
$\operatorname*{\mathcal{U}_{uc}}$$\coloneqq\left\{u_{uc}:\mathbb{R}^{+}\to\mathbb{R}^{p}:\|u_{uc}(t)\|_{\infty}\leq 1~{}\text{for all $t$}\right\}.$ (4b)
In this setting, the system has authority over the controlled input $u_{c}$, but the uncontrolled input $u_{uc}$ can be chosen arbitrarily from the space $\operatorname*{\mathcal{U}_{uc}}$, potentially by an adversary. However, $u_{uc}$ is observable and can be used to design $u_{c}$.
Our objective is the task of finite-time regulation, achieving $x(t_{f})=0$ for a specified final time $t_{f}$ using the control $u\in\operatorname*{\mathcal{U}}$ in the nominal case and $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ for any $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$ in the malfunctioning case. In this context, we define the notion of finite-time stabilizing resilience of a system, adapted from Bouvier et al. (2021).
Definition 1 (Finite-time Stabilizing Resilience)
A system (1) is resilient to the loss of control authority over $p$ of its actuators, represented by the matrix $B_{uc}$ if, for all uncontrolled inputs $u_{uc}\in\operatorname*{\mathcal{U}_{uc}}$ and for a given final time $t_{f}$, there exists a controlled input $u_{c}\in\operatorname*{\mathcal{U}_{c}}$ such that the system achieves finite-time regulation, i.e. $x(t_{f})=0$.
A system may be finite-time stabilizing resilient for some final times $t_{f}$ and not resilient for other final times. Throughout the rest of this paper, we refer to finite-time stabilizing resilience as simply resilience. We also assume that the nominal dynamics (1) are controllable so that finite-time regulation can be achieved from any initial state $x_{0}$. In both the nominal and malfunctioning cases, we are interested in the minimal control energies to achieve this task. Achieving this task might require considerably more control energy in the malfunctioning case compared to the nominal case. We thus aim to quantify the maximal additional control energy used by the malfunctioning system over all possible uncontrolled inputs $u_{uc}$, compared to the nominal system. To this end, we make the following definitions.
Definition 2 (Nominal Energy)
@@
Definition 3 (Malfunctioning Energy)
@@
Definition 4 (Total Energy)
@@
Definition 5 (Worst-case Total Energy)
The worst-case total energy is the maximal effect of the uncontrolled input $u_{uc}$ on the total energy (7) used by the system:
$\overline{E}_{M}(x_{0},t_{f})$ $=\sup_{u_{uc}\in\operatorname*{\mathcal{U}_{uc}}}\left\{E_{M}^{*}(x_{0},t_{f},u_{uc})+\left\|u_{uc}\right\|_{\mathcal{L}_{2}}^{2}\right\}.$ (8)
The goal of this paper is to understand how much larger the worst-case total energy is, compared to the nominal energy. This increase in energy is quantified by the following energetic resilience metric.
Definition 6 (Energetic Resilience)
For an initial condition $x_{0}$ at a distance of at least $R$ from the origin, we define the energetic resilience of system (1) as
@@
Note that if $r_{M}(t_{f},R)\geq 1/2$, then the malfunctioning system uses at most twice the energy used by the nominal system to achieve finite-time regulation, for a given $t_{f}$ and $R$. We thus require a lower bound for this metric. Without the constraint $\|x_{0}\|_{2}\geq R$, $E_{N}^{*}(x_{0},t_{f})$ can be arbitrarily close to zero, resulting in a trivial value for $r_{M}(t_{f},R)$. We note that the definition of $r_{M}(t_{f},R)$ is closely related to the definition of quantitative resilience defined by Bouvier et al. (2021) and Bouvier and Ornik (2023). However, quantitative resilience was defined using reachable time while our definitions consider the control energy.
In the following section, we derive closed-form expressions for the nominal and malfunctioning energy, as well as a bound on the worst-case total energy. To quantify the maximal additional energy used by the malfunctioning system compared to the nominal system, we compare the quantities $E_{N}^{*}(x_{0},t_{f})$ from (5) and $\overline{E}_{M}(x_{0},t_{f})$ from (8). This comparison is achieved in Section 4 by deriving bounds on the energetic resilience metric (9), for the special case of losing authority over a single actuator.
3 Nominal and Malfunctioning Energies
In this section, we derive expressions for the energies defined in Section 2, as well as the corresponding optimal control inputs. We first present a lemma central to the development in this section.
Lemma 1
Let $\mathcal{Z}\coloneqq\Bigl{\{}z:[0,t_{f}]\to\mathbb{R}^{n_{z}}:\frac{1}{t_{f}}\int_{0}^{t_{f}}z(t)\mathrm{d}t=\overline{z}\in\mathbb{R}^{n_{z}}\Bigr{\}}$ be the set of continuous, real vector-valued functions with given mean value $\overline{z}$ in the interval $[0,t_{f}]$. Let $z^{*}(t)$ denote the function with the minimum energy in $\mathcal{Z}$, i.e., $z^{*}\coloneqq\operatorname*{arg\,min}_{z\in\mathcal{Z}}\int_{0}^{t_{f}}z^{T}(t)z(t)\mathrm{d}t.$ Then,
$z^{*}(t)=\overline{z}~{}\text{ for all $t\in[0,t_{f}]$,}$
i.e., the function with minimum energy is a constant.